OpenStudy (sleepyhead314):

In the xy-plane, line t passes through the origin and is perpendicular to the line 4x + y = k, where k is a constant. If the two lines interest at the point (t, t+1), what is the values of t? a) -4/3 b) -5/4 c) 3/4 d) 5/4 e) 4/3

OpenStudy (sleepyhead314):

@mathslover I need help :3 pwease

mathslover (mathslover):

@vishweshshrimali5

OpenStudy (sleepyhead314):

mmm so the equation of the perpendicular line would be some thing like y = (1/4)x + (k+1) ?

OpenStudy (sleepyhead314):

.-. I'm still confused sorry

OpenStudy (vishweshshrimali5):

Its ok. Lets start from the very beginning: (1) Let equation of line be y = mx (2) Since its perpendicular to other line,so, m*(-4) = -1 => m = 1/4 (3) So, eq. becomes y = x/4 => x = 4y (4) Put x = 4y in second eq., we get, 4(4y) + y = k => y = k/(17) So, x = 4y = (4k)/(17) Thus, point of intersection is \((\cfrac{4k}{17}, \cfrac{k}{17})\). But, it should be of the form (t,t+1). So, \(\cfrac{k}{17} - \cfrac{4k}{17} = 1\) \(\implies \cfrac{-3k}{17} = 1\) \(\implies k = \cfrac{-17}{3}\) So, \[t = \cfrac{4k}{17}\] \[\implies t = \cfrac{4}{17}*\cfrac{-17}{3}\] \[\implies t = \cfrac{-4}{3}\]

OpenStudy (vishweshshrimali5):

Did you get this ?

OpenStudy (sleepyhead314):

it's a little confusing... but I got lost at the k/17 - 4k/17 part why did you do that? :/

OpenStudy (vishweshshrimali5):

See the point of intersection was supposed to be (t,t+1) but we got (4k/17, k/17). So on comparing, we get, t = 4k/17 t+1 = k/17 Now, (t+1) - t = 1 => (k/17) - (4k/17) = 1

OpenStudy (sleepyhead314):

ooohhhhhh I see.... right I remember doing something like t + 1 = 1/4 t -> 3t = -4 t = -4/3 but answer key says C

OpenStudy (sleepyhead314):

oh wait nvm

OpenStudy (sleepyhead314):

I got it xD thanks for your help! <3

OpenStudy (vishweshshrimali5):

It was my pleasure friends.